OpenStudy (anonymous):

Center and Radius of x2 + y2 + 10x – 6y + 18 = 0? :)

OpenStudy (accessdenied):

It would be useful to have a form more like \((x - h)^2 + (y - k)^2 = r^2\) here. Are you familiar with completing the square?

OpenStudy (anonymous):

Nope, show me how you'd get this answer step by step? :)

OpenStudy (anonymous):

try to make factors in term of x and y...by addind and substracting constants..

OpenStudy (anonymous):

???

OpenStudy (accessdenied):

The general idea of completing the square is taking a form like this: \(x^2 + ax\) and adding and subtracting something to find a form more like \((x + n)^2 - m\) The process is a little abstract. In general... \((x + n )^2 = x^2 + 2n x + n^2\). If we have \(x^2 + 2nx\), we add an \(n^2\) to simplify that into \((x + n)^2\), and subtract a \(n^2\) to balance out what we just added on so the value is the same as when we started. To change \(2n\) into \(n^2\), we divide by two and then square it. That'll be what we do to the linear coefficient in every case of completing the square.

OpenStudy (waleed_imtiaz):

centre would be (-5,3) do u know how?

OpenStudy (waleed_imtiaz):

and radius would be 4

OpenStudy (anonymous):

How? :)

OpenStudy (accessdenied):

For a real example, which may make that ^ seem simpler... :P We have two quadratics to complete. \(x^2 + 10x\) and \(y^2 - 6y\) "To change \(2n\) into \(n^2\), we divide by two and then square it. That'll be what we do to the linear coefficient in every case of completing the square." For \(x^2 + 10x\), the "2n" is 10. 10/2 = 5. Then we square it, 5^2 = 25. \(x^2 + 10x + 25 - 25\) \((x + 5)^2 - 25\) For \(y^2 - 6y\), its not too different from above, just a small adjustment with the negative. (-6/)/2 = -3, we square it to get (-3)^2 = 9. \(y^2 - 6y + 9 - 9\) Minor adjustment is that \(y^2 - 6y + 9\) factors into \((y - 3)^2\) rather than \((y + 3)^2\). \((y - 3)^2 - 9\)

OpenStudy (waleed_imtiaz):

compare this equation with the general form of the circle... its x^2+y^2+2gx+2fy+c=0 find g and f and c...centre is (-g,-f) and radius is \[\sqrt{g ^{2+f^{2}}-c}\]

OpenStudy (waleed_imtiaz):

now calculate ...its very simple :)

OpenStudy (anonymous):

okay... how do you solve that equation? It has so manyy variables....

OpenStudy (accessdenied):

In the problem... \(x^2 + y^2 + 10x - 6y + 18 = 0\) Pair up the quadratics to complete: \(\color{blue}{x^2 + 10x} + \color{red}{y^2 - 6y} + 18 = 0\) I already did the completing above, so I could just replace the red/blue. Notice, however, that it is the same value because we added and subtracted the same quantity. \(\color{blue}{(x + 5)^2 - 25} + \color{red}{(y - 3)^2 - 9} + 18 = 0\) \(\color{blue}{(x + 5)^2} + \color{red}{(y - 3)^2} - \color{purple}{34} + 18 = 0\) \(\color{blue}{(x + 5)^2} + \color{red}{(y - 3)^2} - 16 = 0\) \((x + 5)^2 + (y - 3)^2 = 16\)

OpenStudy (waleed_imtiaz):

@HopelessMathStudent two variables are here just g and f and c is the constant

OpenStudy (waleed_imtiaz):

simple to solve rather to convert first into standard and then see what is the radius and cntre

OpenStudy (anonymous):

So its 5,-3 and 16? or would iI need to sqrt 16 and get four?

OpenStudy (accessdenied):

yeah, sqrt 16 since the general form is r^2. r^2 = 16 (r > 0), so r = 4

OpenStudy (accessdenied):

and its actually opposite: (-5, 3)

OpenStudy (anonymous):

So its center 5,-3 and radius 4 ? :)

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

in the equation is a positve(the five)

OpenStudy (waleed_imtiaz):

yes.... (-5.3_ and then radius is 4

OpenStudy (accessdenied):

note that the general form is (x - h)^2, and we have (x + 5)^2. If we want (x - h)^2, we have (x - -5)^2

OpenStudy (anonymous):

geez access denied's work looks like a Error on windows lol... i guess that's becausei spread my work out lol)

OpenStudy (anonymous):

or x=-5 x+5=0

OpenStudy (accessdenied):

hmm, do you not see something like this? http://puu.sh/1yBD7 I sometimes worried if it would mess up for others... :P

OpenStudy (anonymous):

that is why x=-5 rather than positive

OpenStudy (anonymous):

okay and the 3 is positive?

OpenStudy (anonymous):

if acces denied says so then yes

OpenStudy (anonymous):

I didn't actually do the problem ahahah

OpenStudy (accessdenied):

yep (y - 3)^2, (y - k)^2 is the general form equivalent. k=3 directly. :)

OpenStudy (anonymous):

Thank you so much. I really appreciate it!

OpenStudy (accessdenied):

I would recommend looking into completing the square & practicing it. It can come up in a lot of places, including here. :P

OpenStudy (accessdenied):

Also, you're welcome! :)

OpenStudy (anonymous):

Got it!